OpenStudy (anonymous):

If lim (x->a) [f(x)g(x)] then the lim is f(a)g(a). Would this be true? If f'(a) exists, then lim (x->a) f(x) = f(a) Would this be false?

OpenStudy (zarkon):

1) consider \(f(x)=x-a\) and \(g(x)=\frac{1}{x-a}\) 2) differentiability implies continuity